OpenStudy (anonymous):

PV=PMT(1-(1+i)^-n)/i PV=13,000 PMT=250 n=? i= 0.002875 25,111= 250 (1-(1+0.002875^-n) /0.002875 I know n= 56.41, but I ant to know the steps it took to get the answer.

OpenStudy (kirbykirby):

\[\text{PV}=\text{PMT}\frac{1-(1+i)^{-n}}{i}\\ 13,000=250\frac{1-(1+0.002875)^{-n}}{0.002875}\\ \, \\ 13,000( 0.002875)=250[1-1.002875^{-n}]\\ \frac{37.375}{250}=1-1.002875^{-n}\\ 0.8505=1.002875^{-n}\\ \log(0.8505)=\log[(1.002875)^{-n}]\\ \log(0.8505)=-n\log(1.002875)\\ \, \\ n=-\frac{\log(0.8505)}{\log(1.002875)}=56.4047\]

OpenStudy (anonymous):

Thank you so much! I've been struggling to solve this problem. :)

OpenStudy (kirbykirby):

:)